Headquarters and Headquarter Company
July 7, 1970
Forward Area: 1-20 Infantry Headquarters is located at LZ Liz and will continued to work with ARVN units in their Area of Operations. At 1345hrs 1-20 Infantry TOC we requested an Urgent Dust Off for 1 Vietnamese child with a cut in the forehead. Charlie Company 26 Engineer picked him up at BS764475 with undetermined cause of the wound, completed from LZ Liz at 1352hrs. At 1745hrs TMF at 1738hrs String 76 was activated by personnel and engaged at 1753hrs with 16 VT rounds of 105 HE Artillery and a Visual Recon Team was requested. At 1750hrs PSYOPS Report at grid BS7548 VIP leaflet drop of 27,000, at grid BS7059 a VIP Leaflet drop of 27,000, at grid BS7460 a VC Chieu Hoi targeted at C 19 was dropped 27,000 and at grid BS7267 a VIP drop of 27,000 was conducted. Also Brigade night PSYOPS was conducted at 1558hrs and targeted at C 19 with a broad cast of 15 minutes and at BS7441 a 15 minute was broadcast targeted at C 219, and Alpha Company 1-20 Infantry turned in ammo and paid out 600 Vietnamese Dollars at BS751571, also their was a Med Cap is conducted daily at Bravo Company 1-20 Infantry at grid BS733653.
Rear Area: Battalion Rear continued their Mission of general support for the battalion and continued their Security Mission for LZ Bronco. SP4 JAMES L. JONES was evac to 27th Surgical Evac Hospital on 7/6/70 (See 7/10/70)
July 7, 1970
At 0805hrs Alpha Company the 16, 18, 36, 38 and 37 elements have closed. At 1510hrs Alpha Company while checking the area of Mechanical Ambush that had been detonated at BS751371, a booby trap was found 15 to 20 feet before the Mechanical Ambush, it was a grenade attached to a trip wire across and obvious route to the Mechanical Ambush. Alpha Company is still moving to their Night Defense Position BS756392.
July 7, 1970
Bravo Company continued their joint Operations with PF and RF units on their Hamlet Pacification and security of Highway 1. At 0645hrs Bravo Company the 68 element is moving to BS760470 and the 18 element is in their Day Logger at BS763572. At 0740hrs Bravo Company 34 element at BS742568 engaged 4 VC who evaded to the south. The VC stopped and engaged the 38 element with sniper fire at 250 meters. Artillery was fired into BS7456 and the enemy returned small arms and M-79 fire. Gun ships checked the area at 0820hrs and the 38 element is sweeping into the area. At 0825hrs the 38 element also picked up 4 detainee 2 males and 2 females at BS741567 and BS749570 and evac to Mo Duc later today. At 1621hrs Bravo Company 18 is taking sniper fire from BS745568 and requesting artillery. There are 3 to 4 enemy 350 meters to their northeast, the present location is at BS742564, small arms fire was returned and area was engaged with artillery with negative results. At 0850hrs Bravo Company the 17 and 36 elements have closed the 18 element is at checkpoint 3 at BS755567 and the 56 Day Logger is at BS735655. At 0850hrs Bravo Company the 18 element has closed, the 38 element released the sharks and are sweeping through the area with M-26 (MACV) blocking. At 0930hrs Bravo Company the 38 element has closed on the Day Logger at BS763572 with the 4 detainees. At 1621hrs Bravo Company 18 element reports they are taking some sniper fire from BS745568 and are now requesting artillery. There are 3 to 4 enemy 350 meters to their northeast. The present location of the 18 element is at BS742564. Small arms fire was returned and the area was engaged with artillery with negative findings still checking the area. At 1900hrs Bravo Company 18 element received sniper fire as they were moving back to their Day Logger from the 38 element. At BS758569 they received 10 to 15 rounds saw 1-3 people, at BS757559 20 to 25 rounds were received and one person was seen. At BS752563 50 rounds with 2 persons spotted when small arms fire was returned and at BS751568 10 to 15 rounds with one person spotted. The area of return fire was checked with negative findings. At 2005hrs Bravo Company the 38 element is moving to their Night Defense Position.
July 7, 1970
At 0900hrs Charlie Company 84 element is moving from their Day Logger to their pick up point. At 0925hrs Charlie Company our present location and Day Logger is at BS783467 and all elements are at the Day Logger. At 1350hrs Charlie Company 54 we request an Urgent Dust Off for 1 US with traumatic amputation of both legs from BS779467. He hit a booby trap, pressure device with no shrapnel and C-4 used as the explosive. He was the 4th pack in crossing a field after changing direction from the east to west then north to south. He was near a hedgerow, line # 30 SP4 Ed Harris. He became a KIA at 23rd Med with multiple frag wounds to his body. At 1900hrs Charlie Company we have moved into their Night Defense Position with multiply ambushes out tonight. SP4 NED HENRY HARRIS LISTED AS KHA.
July 7, 1970
Delta Company continued their Security Mission for LZ Liz. At 1940hrs Delta Company access road ambushes are in their location.
July 7, 1970
Echo Company headquarters and 4.2 Mortar Platoon is location on LZ Liz, where the 4.2 Platoon providing direct fire support for the Battalion. At 0500hrs 4.2 Platoon reports Fire Mission 1 for a total expended of 56 HE rounds. Echo Company Headquarters has the coordination responsibility for Base Defense. At 0726hrs Recon Platoon we are in our Day Logger at BS751456. At 0833hrs the 21 element is in Day Logger at BS753443, the 22 elements in the Day Logger at BS742449 and the 23 is at their Day Logger at BS760454. At 0837hrs the Recon Platoon patrol reports the 22 element at 1500hrs will patrol to checkpoint 1 at BS740458 and checkpoint 2 at BS735454, the 23 element will patrol to checkpoint 1 at BS750465 and checkpoint 2 at BS743465 at 1500hrs. Snakes the 21 will snake at BS745457, the 22 element will snake at BS746466 and the 23 snake will be at BS735655. At 1505hrs Recon Platoon the 22 element is moving to BS740458 and the 23 element is moving to BS750468. At 1620hrs Recon Platoon the 22 element has closed with negative findings and the 21 element is at BS753443. At 2030hrs Recon Platoon the 22 snake is 300 meters north of the preplanned location and the 21 and 23 snakes are in their preplanned location.
Back To Main Page
Back To Contents